OpenStudy (anonymous):

sqrt c^15 is c^7 sqrt c right?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

thank you (:

OpenStudy (jdoe0001):

\(\bf \sqrt{c^{15}} \implies\sqrt{(c^7)^2 \times c} \implies c^7 \sqrt{c}\)

OpenStudy (jdoe0001):

yw